OpenStudy (anonymous):

i need help with this log quetion: Write each as the sum or difference of logs:

OpenStudy (anonymous):

\[\log_{4}(4a^3b^8)\]

OpenStudy (wolfe8):

Note that \[\log a + \log b=\log(ab)\]

OpenStudy (anonymous):

yea, but how can i put that formula into what i have?

OpenStudy (wolfe8):

Well, what do you have in the brackets? Is it not \[4a ^{2}\times b^{8}\]?

OpenStudy (anonymous):

yes. so in reality do i just add those instead?

OpenStudy (wolfe8):

Add the log of each expression, yes. And I believe you can further simplify them if needed.

OpenStudy (anonymous):

so it would be 4a^2b8

OpenStudy (wolfe8):

Here, would this help you? http://www.algebra.com/algebra/homework/logarithm/logarithm.faq.question.202938.html I have to go to class now so I cannot explain immediately. But I will if you still need help when I get home. But your answer is not quite right.

OpenStudy (anonymous):

ah thank you for the little bit of help you did ^.6